I was having trouble solving this question:
 
If the coordinates of a point p be (aCosθ, bSinθ) whereθ is a variable quantity, find locuss of p.
I really dont understand how to solve such problems...

Grade:11

2 Answers

Nicho priyatham
8 years ago
the ans is 
(x/a)2+(y/b)2=1
it is simple u just need to eliminate θ
and for that  
take x=acosθ and y=bsinθ
so cosθ=x/a and sinθ=y/b
now since cos2θ + sin2θ=1 we get
(x/a)2+(y/b)2=1
PLZ APPROVE IF USEFUL incase of any doubts put it in ans box
Jyotinder Singh
8 years ago
Well that was simple! I guess I’ll just get some practice and would get a hang of it. Thanks!